# Iceberg Sensemaking: A Process Model for Critical Data Analysis

### v3

@@ 3.2 Interpretive Schemas
An interpretivist departs from cognitivists and positivists by asserting the variability and contingency of schemas that shape our understanding of things and people in the world. In this domain, schemas may also be called frames [63], sharing with Minsky [51] as well as Goldstein and Papert [52] an interest in delineating which things fall inside and outside a given frame. Crucial to the interpretivist approach to frames is the recognition that any framework may be reframed—challenging its basic assumptions when we find the implications false, unjust, unsound, unfit for the intended purpose, or otherwise lacking. The interpretivist sees empirical data as inseparable from acts of interpretation that occur throughout the process of gathering and analyzing data. Many skeptics of data science readily contend that raw data is an oxymoron concealing the interpretive work that already underlies any dataset from the moment we encounter it [64, 65].
@@
While most early HCI research at Xerox PARC adopted a cognitivist and often positivist lens, as outlined above, the anthropologist Lucy Suchman [66, 67] departed toward a more interpretivist approach in her research at PARC on the notoriously tricky case of photocopier interfaces. Suchman examined users’ interpretive tactics, and especially the breakdown of user understanding, as a result of highly contingent forms of reasoning on the fly while trying to follow instructions for navigating interfaces and other technical tasks. For Suchman, “crucial processes are interactional and circumstantial, located in the relationships among actors and between actors and their embedding situations” (2007, p. 30). Here, Suchman departs from the positivist tradition exemplified by Simon and prevalent among other PARC researchers, finding that the schemas guiding understanding and action are highly personal and contextual[68].
@@
Suchman's work has become something of a classic in the interpretivist domain of Science and Technology Studies (STS), an interdisciplinary field in which historians, sociologists, anthropologists, and philosophers examine the construction of facts and artifacts as consequences of human institutions, values, culture, and politics [69, 70, 71]. Suchman's ethnographic work dovetails with the feminist STS theorist Donna Harraway's [72] influential account of situated knowledge, which treats any claim to knowledge as a positional standpoint of the individual and their context. The STS literature marks a critical departure from the positivist assumption that knowledge can be treated as objective and human-independent, instead emphasizing interpretive complexities, contextual factors, and systems of domination that shape scientific knowledge and technological systems.
@@
Another humanistic model of interpretation that dovetails with the STS literature was presented by Drucker [19] in a recent IEEE VIS capstone as particularly suitable for critical approaches to visualization: hermeneutics. The basic idea behind hermeneutics is that we come to understand the whole of something by examining its parts, and the parts by examining the whole, tracing back and forth in this manner to develop a contextually situated understanding of a subject. This sequential, iterative process of interpretation is often depicted as a hermeneutic circle, although its progressive nature is often likened to a spiral instead. Hermeneutics offers a model of understanding centered on individual observers who may construct different interpretations of the same information due to positional differences and other complexities. While many critical theorists have used hermeneutics to emphasize the ultimate instability of meaning through deconstruction, today this stance is largely outweighed by milder forms of hermeneutics focused on the nature and consequences of interpretation as an act that is always partial, personal, and entangled with contextual factors. Drawing on these productive features of hermeneutic analysis, our model depicts sensemaking as an iterative sequence of interpretations tracing back and forth between schemas and data.
@@
Whereas hermeneutics tends to focus on sensemaking at the individual level, many interpretivists zoom out to the level of social constructionism, emphasizing the role of culture, institutions, and power in the creation of meaning [73]. For a social constructionist, even qualities as basic as color, time, and temperature rest on contingent, subjective schemas for collectively making sense of the world around us. To be a social constructionist does not require denying that these qualities ultimately emerge from physical reality, but rather insists that our subjective understanding precedes and makes possible any pursuit of knowledge. Even data that are gathered, ordered, calculated, and otherwise manipulated by a machine are also socially constructed because the schemas that are durably embedded in these machines originate in social processes and institutional structures with specific histories. This point is the crux of many social constructionist critiques of science and technology, especially in the realm of data. Against the claim that facts can speak for themselves, social constructionists often reply that facts speak for the powerful, whose interests are served by a dominant interpretation expressed in hegemonic appeals to common sense and received wisdom[72, 17]. Critical theorists exemplified by Michel Foucault developed this deeper concern with power by describing the specific ways that modern systems of authority assert control by gathering and manipulating information under the guise of objective knowledge [74, 75]. The power to label, sort, and categorize people and things into schemas becomes calcified in systems that reproduce themselves through disciplinary institutions such as schools, workplaces, prisons, and even mundane systems of classification [76, 77, 78]. Critical accounts of prisons and the construction of criminality have been particularly influential in critiques of data science, where algorithms for predictive policing and parole recommendations make use of schemas that reproduce the systemic racism of the criminal justice system [11]. When a social construction like ‘criminality’ is treated as a basic fact about people, even a perfectly sound sensemaking process will go awry by overlooking its own political entanglements. The algorithmic recommendations given to judges on a parole board may be presented as the calculated likelihood of a specific person committing another crime based on their demographics and past offenses, yet these algorithms are notably skewed in terms of race because racialized minorities are often over-policed systematically, whereas white-collar crime is rarely even tabulated [6, 79].
@@
Beyond philosophical disagreements about the relative soundness of cognitivism and positivism versus interpretivism and social constructionism, data-driven fields as a whole require theoretical foundations capable of identifying and mitigating bias, injustice, and the sort of naive realism that often allows these failures to be treated as the output of an objective scientific process. While many interpretivist critiques of injustice in data science focus on algorithms and predictive systems like machine learning (ML) and artificial intelligence (AI) rather than human sensemaking per se [80, 9, 14], the schemas deployed early in the sensemaking process are often carried over into automated systems, amounting to a form of tacit knowledge that passes unexamined once these systems are up and running [26]. A critical, interpretivist, and social constructionist approach to sensemaking directs attention to unjust, inaccurate, and otherwise misleading schemas at the earliest stages of data analysis and visualization. The process outlined below illustrates one way of modeling how to productively work with data while thinking like an interpretivist.
@@ 4 Model
Figure 4: The Iceberg Sensemaking Model. The blue wavy waterline separates the schemas below from the datasets and findings that emerge from them. Data sources, people, and power are external sources of the schemas that eventually become datasets. Actions are classified into three phases (Add, Check, and Refine); the sensemaking process can loop back to any previous phase before reaching a conclusion.
@@ 4.1 Process
We developed and fine-tuned the model itself through multiple rounds of iteration. In each round, we generated sketches of potential structures, then evaluated their descriptive and prescriptive power through detailed walkthroughs of how these models would apply to several sensemaking scenarios. We developed these scenarios to illustrate the practical application of the model to concrete use cases, and to test the clarity and cogency of the developing model. In some cases, we drew directly from the literature, where we found salient critiques of past sensemaking approaches; in others, we devised an emblematic situation to investigate important implications. For each scenario, we describe both a positive and a negative variant. The positive case illustrates functional data analysis, adhering to the principles of the model; the negative case illustrates dysfunctional sensemaking that this model serves to diagnose. When we noted shortcomings in the evolving model's ability to specify important aspects of these scenario contexts or provide guidance to analysts, we honed the model to sharpen its ability to describe and diagnose the sensemaking process in these practical scenarios. In turn, we also refined and augmented the scenarios as the model developed. In Scenarios (Sec. 5), we present these walkthroughs in detail.
@@ 4.2 Overview
The Iceberg Sensemaking model (Fig. 4) is built around the central metaphor that data is the visible tip of an underlying schematic iceberg. In other words, data always grows out of schemas, not vice versa. The mass of possible schemas always exceeds the dataset given the myriad possibilities of framing data in different ways. We further distinguish between explicit and tacit schemas. The explicit schemas are recorded in some human- or machine-readable way that accompanies what is considered to be the data itself. These explicit schemas can range from annotations, to basic metadata such as column headers, to a complete data dictionary, to a detailed provenance record of how the dataset came to be. Tacit schemas are not explicitly recorded—some are matters of context, some constitute fundamental assumptions in the dataset, and in many cases they are not immediately apparent to the analyst. Tacit schemas include undocumented aspects of the creation and transformation of a dataset, domain knowledge considered common sense by the dataset creators, and unconsidered assumptions that were not surfaced in any previous analysis. If you have ever found the need to stop and examine a previously ignored feature of a dataset that gives it context, you have articulated part of the tacit schema and moved it into the explicit schema. Findings are conclusions that have been evaluated, confirmed, and communicated as the result of the sensemaking process.
@@
In the Add phase, the analyst acquires data, incorporates explicit schemas from the data, and absorbs the tacit schemas of both data and people. At the start of the sensemaking process, an initial dataset is acquired. Acquiring data necessarily results in the incorporation of tacit and explicit schemas underlying it. The absorbed tacit schema of the analyst is their own understanding of the subject matter, which guides them from the beginning of the sensemaking process. The analyst may also absorb the tacit schemas of outside parties, such as managers, colleagues, or other stakeholders. These new schemas may also be specific hypotheses or alternative viewpoints drawn from outside consultation and other information sources such as news coverage of the subject.
@@
In the Refine phase, the analyst should consider the role of power in their work. This consideration may include questions about who made the dataset, people reflected in the data, potential absences in the data, people affected by the results of their analysis, assumptions made at the outset of the project, or even managerial pressure to produce certain findings. The analyst then articulates tacit schemas that were previously unstated. These tacit schemas may arise from any of the three phases: absorbing during the Add phase, evaluating during the Check phase, or considering power in the Refine phase. The articulation process moves previously unstated material into active consideration within the explicit schemas. Articulation of tacit schemas is a choice, and the choices an analyst makes about what to make explicit are fundamentally at the core of sensemaking outcomes. Although a complete articulation of tacit schemas is basically impossible because the full context of both the dataset and the analyst’s mindset may never be known, the more thoroughly we articulate schemas in conversation with other possible schemas, the greater our certainty that we have not taken for granted false, misleading, or even harmful considerations in the background of the sensemaking process. With this newly articulated material, the analyst may use the explicit schema to update the dataset by cleaning, sorting, filtering, recoding, restructuring, and otherwise adapting their datasets. Next, the analyst uses their data to formulate findings. These may be preliminary findings that lead the analyst to loop back and perform another add, check, or refine phase. Eventually the process ends, perhaps due to deadline pressure or the analyst’s subjective sense of satisfaction with the thoroughness of their results. In either case, the final formulation of findings marks the end of the sensemaking process.
Hypothetically, a data analysis scenario could involve only a single pass through the main Add-Check-Refine sequence, but like earlier models, the Iceberg Model encourages the looping nature of the sensemaking process. Analysts typically iterate with many passes through this loop, repeating the cycle of acquiring and incorporating, interpreting and evaluating, considering and articulating and updating, before formulating their final findings. Inner loops may occur at any phase; for example, multiple rounds of interpretation and evaluation could happen before any refinement takes place.
@@ 5.1 Tacit and Explicit Schemas
When sensemaking models use the term schema to mean only the explicitly constructed framework of understanding, they fail to guide analysts to articulate and account for tacit background influences on the sensemaking process. When a sensemaking model deploys the term schema to mean only the explicit schema, it fails to account for knowledge, beliefs, interests, and biases implicit for both the dataset and the analyst, thus neglecting to address and ameliorate possible sources of bias, power, systemic injustice, and even the basic blindspots of different professions and field of study.
Pirolli & Card's [29] treatment of schemas reflects only the creation of the explicit schema, eliding the tacit schemas at play from the beginning of the sensemaking process. The schema plays a larger role in the Grolemund & Wickham model [22], seemingly encompassing both tacit and explicit dimensions in the search for a relevant schema during the exploratory phase and incorporation of new insights during confirmation phase. Nevertheless, because Grolemund & Wickham do not make this distinction, the schema is mainly treated as an explicitly understood model to refine, not an unknown factor to investigate. The Sacha et al. model [23] reflects an overly optimistic treatment of the analyst's starting point as a solid block of reliable knowledge: we prefer to treat the assumptions, models, values, and motivations brought to the sensemaking process as schemas in order to underline that these are constructs in need of evaluation and assessment with respect to alternatives.
@@ 5.2 Schemas First and Always
Although the Pirolli & Card model [29] does discuss schema formation as a framing process, it does not adhere to our Schemas First and Always principle. To treat schemas as devices that enter the sensemaking process midway is to foreground the explicit articulation of a schema, ignoring the tacit schemas that shaped the analyst's thinking. This treatment of schemas places a crucial part of the sensemaking process beyond consideration.
Grolemund & Wickham's [22] cognitive model also depicts schemas entering into sensemaking after the process is already underway. Their depiction of confirmatory analysis could be construed as a schema-first process in which one's initially held schema directs the search for relevant data. Yet in the earlier phase of exploratory analysis, the search for a relevant schema is treated as though it occurs without mental schemas already in play. For a cognitivist who treats schemas as discrete tools that we acquire and place in a kit for later use, it’s easy to imagine starting the sensemaking process with some schemas tucked away, others yet to be acquired, and a basically undecided stance in the absence of a working schema. We contend that schemas are the prime mover of the sensemaking process rather than a gear set in motion at a later stage.
Because Sacha et al. [23] begin and end with knowledge, the element closest to a schema in their model, it might seem to meet our Schemas First and Always principle. The trouble lies in their treatment of knowledge as a simple input defined as “justified belief.” This definition may seem sensible on its face, but philosophers actually formulate this position as “justified true belief” and typically use it as the starting point to argue against the absolute need for truth, justification, and even belief as necessary conditions of knowledge[81, 82]. Even if we approach justification from another angle, not as an essential property but as a social construction residing in collective agreement, this is not a simple property to locate in one belief versus another, but rather something deeply entangled with structural power from the beginning of the sensemaking process. Crucially, what counts as justification may lead to different sensemaking outcomes, so the criteria for justified belief should be treated as a variegated feature of different schemas introduced throughout the sensemaking process. The point is that knowledge is complicated, and it is particularly important for sensemaking models to treat knowledge as complicated.

### v4

@@ 3.2 Interpretive Schemas
An interpretivist departs from cognitivists and positivists by asserting the variability and contingency of schemas that shape our understanding of things and people in the world. In this domain, schemas may also be called frames [64], sharing with Minsky [51] as well as Goldstein and Papert [52] an interest in delineating which things fall inside and outside a given frame. Crucial to the interpretivist approach to frames is the recognition that any framework may be reframed—challenging its basic assumptions when we find the implications false, unjust, unsound, unfit for the intended purpose, or otherwise lacking. The interpretivist sees empirical data as inseparable from acts of interpretation that occur throughout the process of gathering and analyzing data. Many skeptics of data science readily contend that ‘raw data is an oxymoron’, concealing the interpretive work that already underlies any dataset from the moment we encounter it [65, 66].
@@
While most early HCI research at Xerox PARC adopted a cognitivist and often positivist lens, as outlined above, the anthropologist Suchman [67, 68] departed toward a more interpretivist approach in her research at PARC on the notoriously tricky case of photocopier interfaces. Suchman examined users’ interpretive tactics, and especially the breakdown of user understanding, as a result of highly contingent forms of reasoning on the fly while trying to follow instructions for navigating interfaces and other technical tasks. For Suchman, “crucial processes are interactional and circumstantial, located in the relationships among actors and between actors and their embedding situations” (2007, p. 30). Here, Suchman departs from the positivist tradition exemplified by Simon and prevalent among other PARC researchers, finding that the schemas guiding understanding and action are highly personal and contextual[69].
@@
Suchman’s work has become something of a classic in the interpretivist domain of Science and Technology Studies (STS), an interdisciplinary field in which historians, sociologists, anthropologists, and philosophers examine the construction of facts and artifacts as consequences of human institutions, values, culture, and politics [70, 71, 72]. Suchman’s ethnographic work dovetails with the feminist STS theorist Harraway’s [73] influential account of situated knowledge, which treats any claim to knowledge as a positional standpoint of the individual and their context. The STS literature marks a critical departure from the positivist assumption that knowledge can be treated as objective and human-independent, instead emphasizing interpretive complexities, contextual factors, and systems of domination that shape scientific knowledge and technological systems.
@@
Another humanistic model of interpretation that dovetails with the STS literature was presented by Drucker [19] in a recent IEEE VIS capstone as particularly suitable for critical approaches to visualization: hermeneutics. The basic idea behind hermeneutics is that we come to understand the whole of something by examining its parts, and the parts by examining the whole, tracing back and forth in this manner to develop a contextually situated understanding of a subject. This sequential, iterative process of interpretation is often depicted as a hermeneutic circle, although its progressive nature is often compared to a spiral instead. Hermeneutics offers a model of understanding centered on individual observers who may construct different interpretations of the same information due to positional differences and other complexities. While many critical theorists have used hermeneutics to emphasize the ultimate instability of meaning through deconstruction, today this stance is largely outweighed by milder forms of hermeneutics focused on the nature and consequences of interpretation as an act that is always partial, personal, and entangled with contextual factors. Drawing on these productive features of hermeneutic analysis, our model depicts sensemaking as an iterative sequence of interpretations tracing back and forth between schemas and data.
@@
Whereas hermeneutics tends to focus on sensemaking at the individual level, many interpretivists zoom out to the level of social constructionism, emphasizing the role of culture, institutions, and power in the creation of meaning [74]. For a social constructionist, even qualities as basic as color, time, and temperature rest on contingent, subjective schemas for collectively making sense of the world around us. To be a social constructionist does not require denying that these qualities ultimately emerge from physical reality, but rather insists that our subjective understanding precedes and makes possible any pursuit of knowledge. Even data that are gathered, ordered, calculated, and otherwise manipulated by a machine are also socially constructed because the schemas that are durably embedded in these machines originate in social processes and institutional structures with specific histories. This point is the crux of many social constructionist accounts of science and technology, especially in the realm of data. Against the claim that facts can speak for themselves, social constructionists often reply that facts speak for the powerful, whose interests are served by a dominant interpretation expressed in hegemonic appeals to common sense and received wisdom[73, 17]. Critical theorists exemplified by Foucault developed this deeper concern with power by describing the specific ways that modern systems of authority assert control by gathering and manipulating information under the guise of objective knowledge [75, 76]. The power to label, sort, and categorize people and things into schemas becomes calcified in social systems that reproduce themselves through disciplinary institutions such as schools, workplaces, government, prisons, and even everyday systems of classification [77, 78, 79]. In particular, critical accounts of prisons and the construction of criminality have been influential in recent critiques of data science, where algorithms for predictive policing and parole recommendations make use of schemas that reproduce the systemic racism of the criminal justice system [11]. When a social construction like ‘criminality’ is treated as a basic fact about people, even a perfectly sound sensemaking process will go awry by overlooking its own political entanglements. The algorithmic recommendations given to judges on a parole board may be presented as the calculated likelihood of a specific person committing another crime based on their demographics and past offenses, yet these algorithms are notably skewed in terms of race because racialized minorities are often over-policed systematically, whereas white-collar crime is rarely even tabulated [6, 80].
It is important to note that the visualization literature contains an ever-growing thread of work that aligns with interpretivist goals and methods, whether or not this work is labelled with that specific term. For example, the extensively adopted methodology of design studies uses the term ‘qualitative’ rather than ‘interpretivist’ but is explicitly non-positivist, stating that transferability rather than reproducibility is the end goal for all aspects of that work, including schemas constructed by researchers during problem abstraction [81]. Some recent work does invoke interpretivism more explicitly, including deeper considerations of the ethics of collaboration during design studies [82] and particularly larger-scope considerations of epistemological frameworks [20, 83].
@@
Beyond philosophical disagreements about the relative soundness of cognitivism and positivism versus interpretivism and social constructionism, data-driven fields as a whole require theoretical foundations capable of identifying and mitigating bias, injustice, and the sort of naive materialism that allows these failures to be treated as the output of an objective scientific process. While many interpretivist critiques of injustice in data science focus on algorithms and predictive systems like machine learning (ML) and artificial intelligence (AI) rather than human sensemaking per se [84, 9, 14], the schemas deployed early in the sensemaking process are often carried over into automated systems, amounting to a form of tacit knowledge that passes unexamined once these systems are up and running [26]. A critical, interpretivist, and social constructionist approach to sensemaking directs attention to unjust, inaccurate, and otherwise misleading schemas at the earliest stages of data analysis and visualization. The process outlined below illustrates one way of modeling how to productively work with data while thinking like an interpretivist.
@@ 4 Model
Figure 4: The Iceberg Sensemaking Model depicts how a data analyst may use Datasets, Explicit Schemas, and Tacit Schemas to develop Findings through a critical sensemaking process. The wavy waterline separates the Tacit Schemas below from the Explicit Schemas, Datasets, and Findings above. Data Sources and People are external sources of the schemas that eventually become Datasets. The analyst is directed to consider the role of power, in the broad sense of social influence or ideology, to inform their articulation of Tacit Schemas. Actions are classified into three phases (Add, Check, and Refine). The sensemaking process can loop back to any previous phase before reaching a conclusion when Findings are finalized.
@@ 4.1 Process
We developed and fine-tuned the model itself through multiple rounds of iteration. In each round, we generated sketches of potential structures, then evaluated their descriptive and prescriptive power through detailed walkthroughs of how these models would apply to several sensemaking scenarios. We developed these scenarios to illustrate the practical application of the model to concrete use cases, and to test the clarity and cogency of the developing model. In some cases, we drew directly from the literature, where we found salient critiques of past sensemaking approaches; in others, we devised an emblematic situation to investigate important implications. For each scenario, we describe both a positive and a negative variant. The positive case illustrates functional data analysis, adhering to the principles of the model; the negative case illustrates dysfunctional sensemaking that this model serves to diagnose. When we noted shortcomings in the evolving model’s ability to specify important aspects of these scenario contexts or provide guidance to analysts, we honed the model to sharpen its ability to describe and diagnose the sensemaking process in these practical scenarios. In turn, we also refined and augmented the scenarios as the model developed. In Scenarios (Sec. 5), we present these walkthroughs in detail.
@@ 4.2 Overview
The Iceberg Sensemaking model (Fig. 4) is built around the central metaphor that data is the visible tip of an underlying schematic iceberg. In other words, data always grows out of schemas, not vice versa. The mass of possible schemas always exceeds the dataset given the myriad possibilities of framing data in different ways. We further distinguish between explicit and tacit schemas. The explicit schemas are recorded in some human- or machine-readable way that accompanies what is considered to be the data itself. These explicit schemas can range from annotations, to basic metadata such as column headers, to a complete data dictionary, to a detailed provenance record of how the dataset came to be. Tacit schemas are not explicitly recorded—some are matters of context, some constitute fundamental assumptions in the dataset, and in many cases they are not immediately apparent to the analyst. Tacit schemas include undocumented aspects of the creation and transformation of a dataset, domain knowledge considered common sense by the dataset creators, and unconsidered assumptions that were not surfaced in any previous analysis. If you have ever found the need to stop and examine a previously ignored feature of a dataset that gives it context, you have articulated a tacit schema and made it into an explicit schema. Findings are conclusions that have been evaluated, confirmed, and communicated as the final output of the sensemaking process.
@@
In the Add phase, the analyst acquires data, incorporates explicit schemas from the data, and absorbs the tacit schemas of both data and people. At the start of the sensemaking process, an initial dataset is acquired. Acquiring data necessarily results in the incorporation of both explicit schemas and underlying tacit schemas. The absorbed tacit schema of the analyst is their own understanding of the subject matter, which guides them from the beginning of the sensemaking process. The analyst may also absorb the tacit schemas of outside parties, such as managers, colleagues, collaborators, and other stakeholders. These new schemas may also be specific hypotheses or alternative viewpoints drawn from outside consultation and other information sources, such as news coverage and everyday conversation.
@@
In the Refine phase, the analyst should consider the role of power in their work. This consideration may include questions about who made the dataset, people reflected in the data, potential absences in the data, people affected by the results of their analysis, assumptions made at the outset of the project, or even managerial pressure to produce certain findings. The analyst then articulates tacit schemas that were previously unstated. These tacit schemas may arise from any of the three phases: absorbing during the Add phase, evaluating during the Check phase, or considering power in the Refine phase. The articulation process moves previously unstated material into active consideration within the explicit schemas. Articulation of tacit schemas is a choice, and the choices an analyst makes about what to make explicit are fundamentally at the core of sensemaking outcomes. Although a complete articulation of tacit schemas is basically impossible because the full context of both the dataset and the analyst’s mindset may never be known, the more thoroughly we articulate schemas in conversation with other possible schemas, the greater our certainty that we have not taken for granted false, misleading, or even harmful considerations in the background of the sensemaking process. With this newly articulated material, the analyst may use the explicit schema to update the dataset by cleaning, sorting, filtering, recoding, restructuring, and otherwise adapting their datasets. Next, the analyst uses their data to formulate findings. These may be preliminary findings that lead the analyst to loop back and perform another Add, Check, or Refine sequence. Eventually the process ends, perhaps due to deadline pressure or the analyst’s subjective sense of satisfaction with the thoroughness of their results. In either case, the final formulation of findings marks the end of the sensemaking process.
Hypothetically, a data analysis scenario could involve only a single pass through the main Add-Check-Refine sequence, but like earlier sensemaking models, the Iceberg Model encourages looping through the process. Analysts typically iterate with many passes through this loop, repeating the cycle of acquiring and incorporating, interpreting and evaluating, considering and articulating and updating, before formulating their final findings. Inner loops may occur at any phase; for example, multiple rounds of interpretation and evaluation could happen before any refinement takes place.
@@ 5.1 Tacit and Explicit Schemas
When sensemaking models use the term schema to mean only the explicitly constructed framework of understanding, they fail to guide analysts to articulate and account for tacit background influences on the sensemaking process. When a sensemaking model deploys the term schema to mean only the explicit schema, it fails to account for knowledge, beliefs, interests, and biases implicit for both the dataset and the analyst, thus neglecting to address and ameliorate possible sources of bias, power, systemic injustice, and even the basic blindspots of different professions and fields of study.
Pirolli & Card’s [29] treatment of schemas reflects only the creation of the explicit schema, eliding the role of tacit schemas from the beginning of the sensemaking process. The schema plays a larger role in the Grolemund & Wickham model [22], where it seems to encompass both tacit and explicit dimensions in the search for a relevant schema during the exploratory phase, then the incorporation of new insights during the confirmation phase. Nevertheless, because Grolemund & Wickham do not make this distinction, the schema is mainly treated as an explicitly understood model to refine, not an unknown, tacit factor to investigate. The Sacha et al. model [23] reflects an overly optimistic treatment of the analyst’s starting point as a solid block of reliable knowledge. We prefer to treat the assumptions, models, values, and motivations brought to the sensemaking process as schemas in order to underline that these are constructs in need of evaluation and assessment with respect to alternatives.
@@ 5.2 Schemas First and Always
Although the Pirolli & Card model [29] does discuss schema formation as a framing process, it does not adhere to our Schemas First and Always principle. To treat schemas as devices that enter the sensemaking process midway is to foreground the explicit articulation of a schema, ignoring the tacit schemas that shaped the analyst’s thinking. This treatment of schemas places a crucial part of the sensemaking process beyond consideration.
Grolemund & Wickham’s [22] cognitive model also depicts schemas entering into sensemaking after the process is already underway. Their depiction of confirmatory analysis could be construed as a schema-first process in which one’s initially held schema directs the search for relevant data. And yet, in the earlier phase of exploratory analysis, the search for a relevant schema is treated as though it occurs without mental schemas already in play. For a cognitivist who treats schemas as discrete tools that we acquire and place in a kit for later use, it’s easy to imagine starting the sensemaking process with some schemas tucked away, others yet to be acquired, and a basically undecided stance in the absence of a working schema. We contend that schemas are the prime mover of the sensemaking process rather than a gear set in motion at a later stage.
Because Sacha et al. [23] begin and end with knowledge, the element closest to a schema in their model, it might seem to meet our Schemas First and Always principle. The trouble lies in their treatment of knowledge as a simple input defined as “justified belief”. This definition may seem sensible on its face, but philosophers actually formulate this position as “justified true belief” and typically use it as the starting point for elaborate arguments against the absolute need for truth, justification, and even belief as necessary conditions of knowledge[85, 86]. Even if we approach justification from another angle, not as an essential property but as a social construction residing in collective agreement, this is not a simple property to locate in one belief versus another, but rather something deeply entangled with structural power from the beginning of the sensemaking process. Crucially, what counts as justification may lead to different sensemaking outcomes, so the criteria for justified belief should be treated as a variegated feature of different schemas introduced throughout the sensemaking process. The key point is that knowledge is complicated, and it is particularly important for sensemaking models to treat knowledge as complicated.